OpenStudy (anonymous):

Find the values of x for which the series converges, and for these values find its sum. x - x^3 + x^5 - x^7 + x^9 - ...

OpenStudy (sirm3d):

\[\sum_{n=0}^\infty (-1)^n x^{2n+1}=x-x^3+x^5-x^7+\cdots\] use ratio test on the term of the series

OpenStudy (anonymous):

ok thanks ya :D

OpenStudy (john_es):

This serie always converge, right?

OpenStudy (anonymous):

don't know..

OpenStudy (sirm3d):

Ratio Test for \[\sum u_n\] \[\lim_{n\rightarrow \infty}\left|\frac{u_{n+1}}{u_n}\right|<1\] here, \[u_n=(-1)^nx^{2n+1}\] and \[u_{n+1}=(-1)^{n+1}x^{2(n+1)+1}\]

OpenStudy (john_es):

Ok! Not always converge, x must be in some range of values, thanks.

OpenStudy (sirm3d):

\[\left|\frac{u_{n+1}}{u_n}\right|=\left|x^2\right|\] ratio test: \[\lim_{n\rightarrow \infty}|x^2|<1\Rightarrow -1<x<1\]

OpenStudy (john_es):

For the sum, you can fin this usefull, \[S=x−x^3+x^5±…\] \[x^2S=x^3−x^5+x^7±…\] \[S+x^2S=x⇒S=\frac{x}{1+x^2},\ \ \ \ ∀x∈(−1,1)\]